Circular No. 9117

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A. cbatiau@eps.harvard.edu ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET P/2009 WJ_{50} (LA SAGRA) An apparently asteroidal object discovered last November (discovery

observation tabulated below) on CCD images taken remotely by the “La Sagra Sky Survey” (cf. IAUC 9078) was found on Feb. 6 to show cometary appearance on WISE satellite images (cf. IAUC 9115); A. Mainzer, Jet Propulsion Laboratory, reports that the 12- and 22-μm WISE images show an extended coma ∼ 20′′ in diameter with a tail extending at least 35′′ due east in thirteen frames spanning 1.5 days. Following a request by the Central Bureau, M. T. Read obtained three CCD images with the Spacewatch 1.8-m reflector on Feb. 13.2 UT; analysis of Read’s co-added images by J. V. Scotti yields a coma of diameter 6′′ and total magnitude ∼ 22, with a broad fan-shaped extending ∼ 0′.31 and spanning p.a. 41^{o}– 92^{o}. Also, W. H. Ryan and E. V. Ryan obtained CCD images in poor transparency with the McDonald Observatory 2.4-m reflector on Feb. 15.1 that yield red mag 20.1–20.3 with a faint diffuseness toward p.a. ∼ 50^{o}–80^{o}.

2009 UT  α_{2000}  δ_{2000}  Mag.
Nov. 19.87722  3^{h}04^{m}49.^{s}35  +8^{o}15′38′′.8  18.7

The following improved elliptical orbital elements, residuals, and an ephemeris appear on MPEC 2010-D01.

Epoch = 2010 Mar. 25.0 TT

T = 2010 Mar. 12.4032 TT  ω = 27.^{o}1295
e = 0.409774  Ω = 74.9958 }2000.0
q = 1.791561 AU  i = 11.2759

a = 3.035382 AU n^{o} = 0.1863735 P = 5.29 years

COMETS C/2009 Y3 AND C/2009 Y4 (STEREO) Further to IAUC 9057, additional Kreutz sungrazers have been found

on STEREO HI-1 website images (and subsequently on SOHO website images). K. Battams notes that C/2009 Y3 was found by M. Uchina in SOHO C2 website data, where it peaked at mag ∼ 7–7.5 and showed a ‘double nucleus’ that was not visible in HI-1 images. C/2009 Y4 was bright (perhaps near mag 1 at peak), saturating SOHO C3 and C2 images, where the tail was > 1^{o} long; in HI-1 images, the tail was several degrees long.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 Y3 Dec. 19.184 12^{h}50.^{m}3 − 9^{o}58′ HI-1 AW 2010-B28
C/2009 Y4  31.128 13 05.8 −16 31 HI-1 AW 2010-B28
2010 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green